# Enhancing Graph-based Recommendations with Majority-Voting LLM-Rerank Augmentation

### v1

@@ 4.2 Graph Contrastive Learning Paradigm
We continue to employ LightGCN as our graph encoder due to its efficiency and effectiveness, though our framework is compatible with other graph backbones (e.g., NGCF (Wang et al. 2019), XSimGCL (Yu et al. 2023)). Inspired by SimCLR (Chen et al. 2020), we employ a CL paradigm between two representation views derived from the original and augmented data. This approach mitigates distributional shift caused by potential noise introduced during graph augmentation. We share a common embedding matrix $\mathbf{E}^{(0)}$ to encode both the original graph $\mathcal{G}$ and its augmented version $\mathcal{G}_{\text{aug}}$. During each training epoch, node representations are generated as follows:
@@
where $\lambda\in(0,1)$ is a hyperparameter to control weight of CL loss. The main advantages of this CL paradigm are that it leverages LLMs purely for knowledge-guided augmentation without embedding integration, thereby avoiding high-dimensional representations and complex fusion with collaborative signals. Additionally, VoteGCL shares the same learnable embedding matrix for both the original and augmented graphs, enabling more efficient training with reasonable model complexity. A detailed analysis of model complexity is provided in Appendix LABEL:complex.
@@ 5.1 Experimental Setup
Dataset. We evaluate VoteGCL on four real-world datasets: Amazon (Scientific and Book) (Hou et al. 2024a), MovieLens [^1] https://grouplens.org/datasets/movielens/, Netflix (Wei et al. 2024), and Yelp2018 (Wang et al. 2019). The dataset statistics after preprocessing are summarized in Table 2. By default, we select the top-1 result from majority-vote LLM reranking as virtual edges for data augmentation.
Datasets Users Items Inters. Density Aug.
Amazon Scientific 10,995 5,314 76,091 $1.3\times 10^{-3}$ 10,995
Amazon Book 11,000 9,332 212,484 $2.1\times 10^{-4}$ 11,000
Movielen-100K 610 9,701 100,836 $1.7\times 10^{-2}$ 610
Movielen-1M 6000 25,317 914,108 $6.2\times 10^{-3}$ 6000
Yelp2018 11,091 86,646 544,672 $5.3\times 10^{-4}$ 11,091
Netflix 13,187 17,336 68,933 $3.2\times 10^{-4}$ 13,187
Table 2: Dataset statistics
Experimental Setup. We use NDCG@K and Recall@K to evaluate the alignment between recommended items and user preferences, reflecting ranking quality and retrieval accuracy. APLT@K (Abdollahpouri, Burke, and Mobasher 2019) measures the average percentage of long-tail items among the top-K recommendations, serving as an indicator of popularity bias. Following the leave-one-out strategy, we use the latest interaction for testing, the second latest for validation, and the rest for training. Metrics are computed over all items, and average scores are reported. All models are trained for 100 epochs with 256-dimensional embeddings and a learning rate of $1\text{e}^{-3}$; baseline hyperparameters follow their original settings. LightGCN is used as the retrieval model to generate item candidates for LLM reranking. For augmentation, we use the same LLM as prior work gpt-3.5-turbo (Achiam et al. 2023). All experiments are conducted on a single NVIDIA A5000 GPU.
Baselines. We compare VoteGCL against three categories of baseline models: (1) base graph-based recommendation systems, including LightGCN (He et al. 2020), NGCF (Wang et al. 2019), and MF-BPR (Koren, Bell, and Volinsky 2009); (2) traditional augmented graph recommendation methods, such as SGL (Zhou et al. 2023), GraphDA (Fan et al. 2023b), GraphAug (Zhang et al. 2024), and XSimGCL (Yu et al. 2022); and (3) LLM-enhanced recommendation systems, including KAR (Xi et al. 2024), RLMRec (Ren et al. 2024), and LLMRec (Wei et al. 2024). Detailed descriptions of the baselines are provided in Appendix LABEL:experiment.
@@ 5.2 Main results
Dataset Metric MF-BPR LightGCN NGCF GraphAug GraphDA SGL XSimGCL
@@
R@10 0.0712 0.0817 0.0628 0.0643 0.0799 0.0812 0.0857 $0.0659^{4.94\%}$ $0.0901^{5.13\%}$ $\textbf{0.0928}^{13.6\%}$ 8.28%
N@10 0.0365 0.0403 0.0311 0.0268 0.0385 0.0402 0.0416 $0.0375^{20.6\%}$ $0.0428^{2.88\%}$ $\textbf{0.0462}^{14.6\%}$ 11.06%
A@10 0.0015 0.0027 0.0011 0.0056 0.0104 0.0028 0.0098 $0.0061^{454\%}$ $0.0123^{25.5\%}$ $\textbf{0.0145}^{437\%}$ 39.42%
Yelp2018 R@10 0.0304 0.0425 0.0375 0.0352 0.0422 0.0446 0.0435 $0.0401^{6.9\%}$ $0.0442^{1.61\%}$ $\textbf{0.0487}^{14.6\%}$ 8.79%
N@10 0.0147 0.0214 0.0229 0.0125 0.0230 0.0223 0.0221 $0.0201^{0.0\%}$ $0.0234^{10.9\%}$ $\textbf{0.0247}^{15.4\%}$ 7.86%
A@10 0.0075 0.0038 0.0014 0.0012 0.0032 0.0096 0.0068 $0.0018^{28.6\%}$ $0.0210^{208\%}$ $\textbf{0.0349}^{818\%}$ 263.54%
@@
R@10 0.1173 0.1251 0.0941 0.0986 0.1204 0.1297 0.1291 $0.1317^{39.9\%}$ $0.1335^{3.4\%}$ $\textbf{0.1385}^{10.7\%}$ 6.78%
N@10 0.0850 0.0853 0.0648 0.0646 0.0758 0.0911 0.0896 $0.1008^{48.0\%}$ $0.0936^{4.5\%}$ $\textbf{0.0993}^{16.4\%}$ 9.01%
A@10 0.2563 0.2015 0.1839 0.0413 0.2034 0.2504 0.1779 $0.2542^{38.2\%}$ $0.1975^{11.0\%}$ $\textbf{0.2795}^{38.7\%}$ 9.05%
Table 3: Performance Comparison of Graph Recommendation Methods on Multiple Datasets Before and After Data Augmentation. Each value in augmented data is shown as $\text{value}^{\text{improvement}}$, where the superscript indicates the percentage improvement over the baseline. Underlined values highlight the best baseline performance, while bold values highlight the best performance after data augmentation. ”Improv” refers to the absolute improvement between these two.
Figure 5: Performance of LLM-enhanced graph recommenders on three datasets under @10 and @20 settings, using dataset-specific models and leveraging available LLM-augmented data for fair comparison.
Traditional graph recommendation
[…]
LLM-enhanced graph recommendation system
As shown in Figure 5, LightGCN + VoteGCL consistently outperforms all other methods across three benchmark datasets (Netflix, Amazon Book, and Yelp2018) under both @10 and @20 evaluation settings, achieving superior performance in terms of both accuracy and diversity. This includes outperforming traditional baselines (LightGCN, SGL) as well as LLM-enhanced models. While LLMRec, RLMRec, and KAR demonstrate notable improvements in ranking metrics such as NDCG particularly when textual embeddings from LLMs are incorporated into the recommendation backbone, they also introduce a higher degree of popularity bias. This drawback is reflected in their lower APLT scores compared to SGL, which applies conventional data augmentation without relying on LLM-generated semantics. In contrast, VoteGCL alleviates this issue by generating high-confidence synthetic interactions through LLM prompting and integrating them effectively via graph contrastive learning, resulting in a more balanced trade-off between recommendation accuracy and fairness.
@@ 5.3 Ablation Studies
Number of Virtual Edges for Augmentation
[…]
Study of VoteGCL
Figure 7: Performance of VoteGCL under different contrastive weight
@@
Table 5 presents an ablation study evaluating the impact of key components in VoteGCL across three datasets. The full VoteGCL model consistently outperforms all baselines in both NDCG@10 and APLT@10. When the graph CL module is removed (”w/o GCL”), performance drops across all datasets, though it still outperforms LightGCN in certain metrics. This suggests that while the LLM-based augmentation is beneficial, without GCL the model becomes more sensitive to noise from the synthetic data. Additionally, removing the LLM reranking with majority voting (”w/o LLM”) and selecting top-1 candidates without using LLM to vote leads to a significant performance collapse. This highlights the importance of majority voting in stabilizing the augmentation and leveraging textual descriptions modalities, making both GCL and LLM-based reranking essential for the effectiveness of VoteGCL.
Zero-shot vs Few-shot Prompting LLM with Different Number of Majority-voting
Figure 6 illustrates the impact of prompting strategy and the number of reranking rounds on the NDCG@10 performance of LLM-based item reranking. Across both datasets, few-shot prompting consistently outperforms zero-shot prompting, highlighting its effectiveness in capturing user preferences through collaborative signals. As the number of majority-voting rounds increases, performance improves and becomes more stable, particularly from 8 rounds onward, for both prompting strategies. Notably, the few-shot setting yields higher mean scores across 30 trials, indicating greater reliability and robustness in reranking outcomes. These findings suggest that effectively leveraging the LLM’s parametric knowledge via few-shot prompting leads to higher-quality augmented data and ultimately enhances the performance of VoteGCL. Notably, the performance at 6 voting rounds is already competitive, making it a practical and cost-effective choice in resource-constrained scenarios. Based on these observations, we recommend combining few-shot prompting with a sufficiently large number of majority-voting rounds (e.g., 6$\geq$ 6≥ 6) to fully realize the benefits of LLM-based reranking in data augmentation.
Controling Hyperparam $\lambda$ Contrastive Weight
As illustrated in Figure 7, the contrastive loss weight $\lambda\in(0,1)$ plays a critical role in balancing recommendation accuracy and popularity bias mitigation. When $\lambda$ increases, the APLT@10 metric improves significantly, indicating greater inclusion of long-tail items and enhanced diversity in recommendations. This effect is consistent across both MovieLens-100K and Netflix datasets. However, this gain in diversity comes with a trade-off NDCG@10 gradually declines as $\lambda$ grows, reflecting a reduction in ranking precision. The results suggest that overly high values of $\lambda$ (e.g., 0.2) may overly prioritize long-tail exploration at the expense of accuracy. Based on these observations, we recommend tuning $\lambda$ within the range of 0.01 to 0.1 to strike an effective balance between accuracy and fairness, depending on the characteristics of the target dataset.
Additional Experimental Results
Additional ablation studies are provided in Appendix LABEL:experiment. These include evaluations of different retrieval models for generating candidate items for LLM reranking, assessments of alternative LLMs for data augmentation (e.g., Qwen2.5 (Yang et al. 2024)), and an analysis of how varying the number of top-K items in the majority-voting process affects performance. Furthermore, we compare the cost of generating augmented data using VoteGCL against other LLM-based methods, demonstrating that our approach offers a cost-efficient solution.

### v3

@@ 4.2 Graph Contrastive Learning Paradigm
We continue to employ LightGCN as our graph encoder due to its efficiency and effectiveness, though our framework is compatible with other graph backbones (e.g., NGCF (Wang et al. 2019), XSimGCL (Yu et al. 2023)). Inspired by SimCLR (Chen et al. 2020), we employ a CL paradigm between two representation views derived from the original and augmented data. This approach mitigates the distributional shift caused by potential noise introduced during graph augmentation. We share a common embedding matrix $\mathbf{E}^{(0)}$ to encode both the original graph $\mathcal{G}$ and its augmented version $\mathcal{G}_{\text{aug}}$. During each training epoch, node representations are generated as follows:
@@
where $\lambda\in(0,1)$ is a hyperparameter to control the weight of CL loss. The main advantages of this CL paradigm are that it leverages LLMs purely for knowledge-guided augmentation without embedding integration, thereby avoiding high-dimensional representations and complex fusion with collaborative signals. Additionally, VoteGCL shares the same learnable embedding matrix for both the original and augmented graphs, enabling more efficient training with reasonable model complexity. A detailed analysis of model complexity is provided in Appendix B.
@@ 5.1 Experimental Setup
Datasets. We evaluate VoteGCL on four real-world datasets from diverse domains: Amazon (Scientific and Book) (Hou et al. 2024a), MovieLens (Harper and Konstan 2015), Netflix (Wei et al. 2024), and Yelp2018 (Wang et al. 2019). Dataset statistics after preprocessing are summarized in Table 5. By default, we select the top-1 result from majority-vote LLM reranking as virtual edges for data augmentation.
Metrics. We use NDCG@K and Recall@K to evaluate alignment between recommended items and user preferences, and APLT@K (Abdollahpouri, Burke, and Mobasher 2019) to measure the average percentage of long-tail items among top-K recommendations as an indicator of popularity bias. Following the leave-one-out strategy, we use the latest interaction for testing, the second latest for validation, and the remaining for training. LightGCN generates item candidates for LLM reranking using gpt-3.5-turbo (Achiam et al. 2023) as prior work.
Baselines. We compare VoteGCL against three categories: (1) base graph-based recommendation systems including LightGCN (He et al. 2020), NGCF (Wang et al. 2019), and MF-BPR (Koren, Bell, and Volinsky 2009); (2) traditional augmented graph recommendation methods such as SGL (Zhou et al. 2023), GraphDA (Fan et al. 2023b), GraphAug (Zhang et al. 2024), and XSimGCL (Yu et al. 2022); and (3) LLM-enhanced recommendation systems including KAR (Xi et al. 2024), RLMRec (Ren et al. 2024), and LLMRec (Wei et al. 2024). We provide details of baselines in Appendix A.
@@ 5.2 Main Results
Dataset Metric
MF-
BPR
Light
GCN
NGCF
Graph
Aug
Graph
DA
SGL XSimGCL
@@
[…]
@@
R@10 0.1173 0.1251 0.0941 0.0986 0.1204 0.1297 0.1291 $0.1317^{39.9\%}$ $0.1335^{3.4\%}$ $\textbf{0.1385}^{10.7\%}$ 6.8%
N@10 0.0850 0.0806 0.0648 0.0646 0.0758 0.0911 0.0896 $\textbf{0.1008}^{48.0\%}$ $0936^{4.5\%} 0.0$$0993^{23.2\%}$ 10.6%
A@10 0.2563 0.2015 0.1839 0.0413 0.2034 0.2504 0.1779 $0.2542^{38.2\%}$ $0.1975^{11.0\%}$ $\textbf{0.2795}^{38.7\%}$ 9.1%
Table 2: Performance Comparison of Vanilla Graph Recommendation Methods on Multiple Datasets Before and After Data Augmentation. Each value in augmented data is shown as $\text{value}^{\text{improvement}}$, where the superscript indicates the percentage improvement over the baseline. Underlined values highlight the best baseline performance, while bold values highlight the best performance after data augmentation. ”Improv” refers to the absolute improvement between these two.
Figure 4: Performance of LLM-enhanced graph recommenders on three datasets under Top K settings (K = 10 and 20), using dataset-specific models and leveraging available LLM-augmented data for fair comparison.
Traditional graph recommendation.
Table 2 demonstrates that the proposed VoteGCL framework exhibits strong model-agnostic capabilities, consistently improving performance across diverse graph-based recommendation architectures, including LightGCN, XSimGCL, and NGCF. VoteGCL achieves substantial improvements in recommendation accuracy, with LightGCN+VoteGCL showing the most significant gains of $13.6\%$ on MovieLens 1M, and $14.6\%$ on Yelp2018 for R@10 metrics. Notably, the framework demonstrates exceptional scalability, delivering consistent improvements across both large-scale datasets (Yelp2018 and MovieLens 1M) and smaller datasets (Amazon Scientific and MovieLens 100K), with LightGCN+VoteGCL achieving particularly strong performance gains of $43.6\%$ on MovieLens 100K and $16.4\%$ on Amazon Scientific for NDCG@10, demonstrating superior adaptation to user preferences in reranking scenarios. More importantly, the framework effectively addresses popularity bias by dramatically enhancing long-tail item recommendations, as evidenced by APLT improvements ranging from $25.5\%$ to $818\%$ across different dataset-model combinations. These results demonstrate that VoteGCL successfully mitigates popularity bias while maintaining recommendation accuracy, promoting diversity by enabling models to learn better representations for underrepresented items regardless of dataset scale. The consistent performance gains across four distinct datasets spanning different domains and varying sizes validate the robust generalizability and universal applicability of VoteGCL.
LLM-enhanced graph recommendation system.
As shown in Figure 4, LightGCN + VoteGCL consistently outperforms all other methods across three benchmark datasets (Netflix, Amazon Book, and Yelp2018) under both @10 and @20 evaluation settings, achieving superior performance in terms of both accuracy and diversity. This includes outperforming traditional baselines (LightGCN, SGL) as well as LLM-enhanced models. While LLMRec, RLMRec, and KAR demonstrate notable improvements in ranking metrics such as NDCG, particularly when textual embeddings from LLMs are incorporated into the recommendation backbone, they also introduce a higher degree of popularity bias. This drawback is reflected in their lower APLT scores compared to SGL, which applies conventional data augmentation without relying on LLM-generated semantics. In contrast, VoteGCL alleviates this issue by generating high-confidence synthetic interactions through LLM prompting and integrating them effectively via graph contrastive learning, resulting in a more balanced trade-off between recommendation accuracy and fairness.
@@ 5.3 Ablation Studies
Figure 6: Performance of VoteGCL under different contrastive weights
@@
Number of Virtual Edges for Augmentation.
Table 3 examines the impact of augmenting different user quantiles on VoteGCL’s performance across three datasets. The results show that expanding augmentation to a larger portion of users does not consistently improve performance in either NDCG@10 or APLT@10. In some cases, performance even declines, suggesting that indiscriminate augmentation can introduce noise that offsets the benefits of additional interactions. These findings highlight the importance of targeted augmentation, specifically, focusing on the bottom $25\%$ of users with the fewest interactions. This approach improves generalization for cold-start users while maintaining overall model robustness and computational efficiency. Alternatively, the trade-off between recommendation accuracy and diversity can be more effectively managed by tuning the contrastive loss weight $\lambda$, allowing finer control over model behavior without increasing augmentation overhead.
Zero-shot vs Few-shot Prompting LLM with Different Number of Majority-voting.
Figure 5 illustrates the impact of the prompting strategy and the number of reranking rounds on the NDCG@10 performance of LLM-based item reranking. Across both datasets, few-shot prompting consistently outperforms zero-shot prompting, highlighting its effectiveness in capturing user preferences through collaborative signals. As the number of majority-voting rounds increases, performance improves and becomes more stable, particularly from 8 rounds onward, for both prompting strategies. Notably, the few-shot setting yields higher mean scores across 30 trials, indicating greater reliability and robustness in reranking outcomes. These findings suggest that effectively leveraging the LLM’s parametric knowledge via few-shot prompting leads to higher-quality augmented data and ultimately enhances the performance of VoteGCL. The performance at 6 voting rounds is competitive, making it a practical and cost-effective choice in resource-constrained scenarios. From these observations, we recommend combining few-shot prompting with a sufficiently large number of majority-voting rounds (e.g., $\geq 6$) to fully realize the benefits of LLM-based reranking in data augmentation.
Controling Hyperparam $\lambda$ Contrastive Weight.
As illustrated in Figure 6, the contrastive loss weight $\lambda\in(0,1)$ plays a critical role in balancing recommendation accuracy and popularity bias mitigation. When $\lambda$ increases, the APLT@10 metric improves significantly, indicating greater inclusion of long-tail items and enhanced diversity in recommendations. This effect is consistent across both MovieLens-100K and Netflix datasets. However, this gain in diversity comes with a trade-off: NDCG@10 gradually declines as $\lambda$ grows, reflecting a reduction in ranking precision. The results suggest that overly high values of $\lambda$ (e.g., 0.2) may overly prioritize long-tail exploration at the expense of accuracy. Based on these observations, we recommend tuning $\lambda$ within the range of 0.01 to 0.1 to strike an effective balance between accuracy and fairness, depending on the characteristics of the target dataset.
Study of VoteGCL.
@@
Table 4 presents an ablation study evaluating the impact of key components in VoteGCL across three datasets. The full VoteGCL model consistently outperforms all baselines in both NDCG@10 and APLT@10. When the graph CL module is removed (”w/o GCL”), performance drops across all datasets, though it still outperforms LightGCN in certain metrics. This suggests that while the LLM-based augmentation is beneficial, without GCL, the model becomes more sensitive to noise from the synthetic data. Additionally, removing the LLM reranking with majority voting (”w/o LLM augment.”) and selecting top-1 candidates without using LLM to vote leads to a significant performance collapse. This highlights the importance of majority voting in stabilizing the augmentation and leveraging textual descriptions modalities, making both GCL and LLM-based reranking essential for the overall robustness and effectiveness of VoteGCL.
Additional Experimental Results.
Additional ablation studies are provided in Appendix A. These include evaluations of different retrieval models for generating candidate items for LLM reranking, assessments of alternative LLMs for data augmentation (e.g., Qwen2.5 (Yang et al. 2024)), and an analysis of how varying the number of top-K items in the majority-voting process affects performance. Furthermore, we compare the cost of generating augmented data using VoteGCL in terms of price per accuracy gain and total generation cost, demonstrating that our approach provides a more cost-efficient alternative to other LLM-based methods.